OpenStudy (anonymous):

1. Check all correct answers (there may be more than one). In the expression a^b, A. The value b is an exponent. B. The value a tells you how many times to multiply b by itself. C. The value a must be > 1. D. The value b tells you how many times to multiply a by itself. E. The value b must be an integer.

OpenStudy (anonymous):

a , d

OpenStudy (anonymous):

and e

OpenStudy (anonymous):

thanks for helping

OpenStudy (solomonzelman):

ONLY a and d @wendylisette, @Rpayne91 Should I explain?

OpenStudy (anonymous):

Yes. Please explain.

OpenStudy (solomonzelman):

OK, you know that b is exponent right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@SolomonZelman its not always an integer ?

OpenStudy (solomonzelman):

\[i t~~can~~be,~~~~~~~a ^{3/4}~~~which~~is,~~~\sqrt[4]{a^3}\]

OpenStudy (solomonzelman):

So no, it doesn't HAVE to be an integer.

OpenStudy (anonymous):

oh ok i see now , thanks

OpenStudy (solomonzelman):

Also \[2^3= 2 \times 2 \times 2\]So b tells us how many times do we multiply a by itself, so it's d, and not b.

OpenStudy (anonymous):

Thanks for the example. I understand now.